water is ejected from a hose of radius 2.5 cm at a speed of 8 m/s. what is the volume of flow rate through the hose?
1
Expert's answer
2020-06-29T14:04:04-0400

We can find the volume of flow rate through the hose from the formula:


"Q = Av,"

here, "A = \pi r^2" is the cross-sectional area of the hose, "v" is the speed of the water.

Then, we get:


"Q = \pi r^2v,""Q = \pi \cdot (0.025 \ m)^2 \cdot 8 \dfrac{m}{s} = 0.016 \ \dfrac{m^3}{s}."

Answer:

"Q = 0.016 \ \dfrac{m^3}{s}."